# Position Paper: What Can Large Language Models Tell Us about Time Series Analysis

### v1

@@ B.1 Overview of Related Works
In the realm of leveraging LLMs as agents for general-purpose time series analysis is still nascent. In the following, we provide an overview of related approaches across different modalities, focusing on strategies for developing robust, general-purpose time series agents. These methods fall into two primary categories. (1) External knowledge integration: this strategy employs ICL prompts to enhance LLMs’ understanding of specific domains. Yang et al. embeds object descriptions and relationships into prompts to aid LLMs in image query analysis (Yang et al., 2022b). Similarly, Da et al. uses prompts containing traffic states, weather types, and road types for domain-informed inferences (Da et al., 2023a). Other studies like (Huang et al., 2022; Singh et al., 2023) include state, object lists, and actions in prompts, allowing LLMs to plan across varied environments and tasks. Wu et al. introduces a prompt manager for ChatGPT to leverage pretrained vision models (Wu et al., 2023), while SocioDojo (Anonymous, 2024a) employs ICL for accessing external knowledge sources like news and journals for decision-making. Despite their efficiency and no need for additional training, these prompt-based methods face limitations such as input length constraints and difficulties in capturing complex time series patterns linguistically. (2) Alignment of LLMs to target modality content: this method aligns LLMs with specific modality content. Schick et al. enables LLMs to annotate datasets with API calls, fine-tuning them for diverse tool usage (Schick et al., 2023). LLaVA (Liu et al., 2023b) generates multimodal language-image instruction data using GPT-4, while Pixiu (Xie et al., 2023) creates a multi-task instruction dataset for financial applications, leading to the development of FinMA, a financial LLM fine-tuned for various financial tasks. Yin et al. offers a multi-modal instruction tuning dataset for 2D and 3D understanding, helping LLMs bridge the gap between word prediction and user instructions (Yin et al., 2023). However, designing comprehensive instructions remains a complex task (Zhang et al., 2023d), and there’s concern that this approach may favor tasks over-represented in the training data (Gudibande et al., 2023).
@@ B.2 Demonstrations
Sit
Instances where there is relatively low movement and consistent values in the accelerometer and gyroscope readings, typical of a sedentary position.
Stand
Instances where there is minimal movement, but the sensor readings may show more variability compared to sitting. Standing typically involves slight variations in body position and may exhibit more fluctuations in sensor readings.

### v2

@@ B.1 Overview of Related Works
In the realm of leveraging LLMs as agents for general-purpose time series analysis is still nascent. In the following, we provide an overview of related approaches across different modalities, focusing on strategies for developing robust, general-purpose time series agents. These methods fall into two primary categories. (1) External knowledge integration: this strategy employs ICL prompts to enhance LLMs’ understanding of specific domains. Yang et al. embeds object descriptions and relationships into prompts to aid LLMs in image query analysis (Yang et al., 2022b). Similarly, Da et al. uses prompts containing traffic states, weather types, and road types for domain-informed inferences (Da et al., 2023). Other studies like (Huang et al., 2022; Singh et al., 2023) include state, object lists, and actions in prompts, allowing LLMs to plan across varied environments and tasks. Wu et al. introduces a prompt manager for ChatGPT to leverage pretrained vision models (Wu et al., 2023), while SocioDojo (Cheng & Chin, 2024) employs ICL for accessing external knowledge sources like news and journals for decision-making. Despite their efficiency and no need for additional training, these prompt-based methods face limitations such as input length constraints and difficulties in capturing complex time series patterns linguistically. (2) Alignment of LLMs to target modality content: this method aligns LLMs with specific modality content. Schick et al. enables LLMs to annotate datasets with API calls, fine-tuning them for diverse tool usage (Schick et al., 2024). LLaVA (Liu et al., 2023b) generates multimodal language-image instruction data using GPT-4, while Pixiu (Xie et al., 2023) creates a multi-task instruction dataset for financial applications, leading to the development of FinMA, a financial LLM fine-tuned for various financial tasks. Yin et al. offers a multi-modal instruction tuning dataset for 2D and 3D understanding, helping LLMs bridge the gap between word prediction and user instructions (Yin et al., 2024). However, designing comprehensive instructions remains a complex task (Zhang et al., 2023b), and there’s concern that this approach may favor tasks over-represented in the training data (Gudibande et al., 2023).
@@ B.2 Demonstrations
Sit Instances where there is relatively low movement and consistent values in the accelerometer and gyroscope readings, typical of a sedentary position.
Stand Instances where there is minimal movement, but the sensor readings may show more variability compared to sitting. Standing typically involves slight variations in body position and may exhibit more fluctuations in sensor readings.